Find the following.
f(x) = 5x − 5, g(x) = 4 − x
(a) (f + g)(x) =

(b) (f − g)(x) =

(c) (fg)(x) =

(d) (f/g)(x) =
What is the domain of
f/g?
(Enter your answer using interval notation.)

Answer by Alan3354(69443)

Find the following.
f(x) = 5x − 5, g(x) = 4 − x
(a) (f + g)(x) =
Add them
===================
(b) (f − g)(x) =
Subtract g(x) from f(x)
========================
(c) (fg)(x) =
Multiply the,
==========================
(d) (f/g)(x) =
Divide f(x) by g(x)
===============
What is the domain of
f/g?